George and Lisa would like to build a double garage in their backyard. They estimate it will cost $15,000 and they would like to pay for it without using financing. If they are willing to wait for 4 years, how much do they need to invest now in order to have $15,000 if they can receive simple interest of 5% per year?
Answer by stanbon(75887)

George and Lisa would like to build a double garage in their backyard. They estimate it will cost $15,000 and they would like to pay for it without using financing. If they are willing to wait for 4 years, how much do they need to invest now in order to have $15,000 if they can receive simple interest of 5% per year?
-------------------
A(t) = P(1 + rt)
15000 = P(1 + 0.05*4)
15000 = P(1.2)
P = $12,500
===================
Cheers,
Stan H.